Four forces of magnitude P,2P,3P and 4P act aling the four sides of square ABCD in cyclic order. Use the vector method to find the resultant force?

Grade:11

2 Answers

Rahul Parkar
14 years ago

side 1 parallel side 3

considering this as X-axis,

P-3P=-2P i          (→ +ive)

for Y-axis

2P-4P=-2P j        (↑ +ive)

mod(∑F)=√(Fx2+Fy2)

            =√(4P2+4P2)

            =2P√2

 

Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
We have,
FAB = P i
FBC = 2P j
FCD = 3P ( – i)
FDA = 4P ( – j)
Hence, resultant vector,
Fr = FAB + FBC + FCD + FDA
    = (P i – 3P i) + (2P j – 4P j)
    = – 2P i – 2P j
Hence, |Fr| = [(-2P)2 + (-2P)2]1/2
Hence, |Fr| = 2√2 P
 
Thanks and regards,
Kushagra